OpenStudy (anonymous):

A guy wire from the top of a tower to a point on level ground 300.0 feet from its base makes an angle of 64 degrees with the ground. If another wire runs to a point 240.0 feet from the base. what angle does it make with the ground?

OpenStudy (amistre64):

tan(64)= h/300 h = 300tan(64) new angle = tan-1(h/240) new angle = tan-1(300tan(64)/240)

OpenStudy (amistre64):

i get 68.6849......

OpenStudy (anonymous):

Yep rounded off it is 68.7